An express train and a local train leave Gray’s Lake at 3 P.M. and head for Chicago 50 miles away. The express travels twice as fast as the local and arrives 1 hour ahead of the local. Find the speed of each train.
Answer by checkley71(8403)

DISTANCE=RATE*TIME
FOR THE EXPRESS TRAIN 50=2X(T-1)
FOR THE LOCAL TRAIN 50=X*T
2XT-2X=XT
2XT-XT-2X0
XT-2X=0
X(T-2)=0
T-2=0
T=2 HOURS FOR THE LOCAL TRAIN.
50=2X
X=25 MPH FOR THE LOCAL TRAIN.
50=2*25(2-1)
2-1=1 HOUR FOR THE EXPRESS TRAIN.
50=2*25*1
50=50